Haver Company currently produces component RX5 for its sole product. The current cost per unit to manufacture the required 50Incremental cost of buying RX5 Buying the units Total incremental cost of buying 50,000 unitsShould the company make or buy RX5?
Answer #1
1
Incremental cost of making RX5 Making the units
Total direct materials 250000 =50000*5
Total direct labor 400000 =50000*8
Variable overhead costs 90000 =50000*9*20%
Total Incremental cost of making 50000 units 740000
2
Incremental cost of buying RX5 Buying the units
Cost to buy the units 900000 =50000*18
Total Incremental cost of buying 50000 units 900000
3
Make the units